# New versions of Frobenius and integral closure of ideals

### v2

@@ 3. The closure operations
In this section we define four new closure operations whose origin is a formal adaptation of the tight closure adjustments of Hara-Yoshida [HY03] and Vraciu [Vra08] (see 3.9) to integral and Frobenius closures. We also note that in this section we use the word “closure” loosely, in that the Hara-Yoshida versions of the closures are not generally idempotent operations, so that the Hara-Yoshida “closures” may not be closed.
@@ 3.1. The $\mathfrak{a}^{t}$-integral closures
In this subsection, $R$ is a ring of arbitrary characteristic, and we will assume the ideals we denote $I$ and $\mathfrak{a}$ intersect $R^{\circ}$ non-trivially. We begin with the adjustments to the ordinary integral closure of an ideal $I$.
@@
Let $I,\mathfrak{a}\subset R$ be ideals and let $t\in\mathbb{R}_{>0}$. Then, the Hara-Yoshida $\mathfrak{a}^{t}$-Integral closure of $I$, denoted $\overline{I}^{\mathfrak{a}^{t}}$, is defined by
$\overline{I}^{\mathfrak{a}^{t}}=\left\{x\in R\left|\text{ for some }c\in R^{\circ}\text{ and all }n\gg 0,\,cx^{n}\mathfrak{a}^{\lceil tn\rceil}\subset I^{n}\right.\right\}.$
Similarly, the Vraciu $\mathfrak{a}^{t}$-Integral closure of $I$, denoted ${}^{\mathfrak{a}^{t}}\overline{I}$, is defined by
${}^{\mathfrak{a}^{t}}\overline{I}=\left\{x\in R\left|\text{ for some }c\in R^{\circ}\text{ and all }n\gg 0,\,cx^{n}\mathfrak{a}^{\lceil tn\rceil}\subset\mathfrak{a}^{\lceil tn\rceil}I^{n}\right.\right\}.$
We first note that the “for all $n\gg 0$” in the definitions above may be strengthened to “for all $n$”, so long as $I$ is of positive height.
@@
Let $I,\mathfrak{a}\subset R$ be ideals and let $t\in\mathbb{R}_{>0}$. Then $x\in\overline{I}^{\mathfrak{a}^{t}}$ (respectively $x\in{}^{\mathfrak{a}^{t}}\overline{I}$) if and only if there exists $c\in R^{\circ}$ such that, for all $n\geq 0$, $cx^{n}\mathfrak{a}^{\lceil tn\rceil}\subset I^{n}$ (respectively $cx^{n}\mathfrak{a}^{\lceil tn\rceil}\subset I^{n}\mathfrak{a}^{\lceil tn\rceil}$).
@@
The proofs of both parts are similar, so we provide the proof for the Hara-Yoshida version only. Let $x\in\overline{I}^{\mathfrak{a}^{t}}.$ Then there exist $n_{0}\in\mathbb{N}$ and $c\in R^{\circ}$ such that $cx^{n}\mathfrak{a}^{\lceil tn\rceil}\subset I^{n}$ for all $n\geq n_{0}.$ Choose $b\in R^{\circ}\cap I^{n_{0}}$. Then $bcx^{n}\mathfrak{a}^{\lceil tn\rceil}\subset I^{n}$ for all $n\geq 0.$
We now record several basic properties of these closure operations. Notably, as long as $I$ and $\mathfrak{a}$ avoid the minimal primes of $R$, the Vraciu $\mathfrak{a}^{t}$-integral closure agrees with the usual integral closure, but the Hara-Yoshida version is generally larger.
@@
Let $I,\mathfrak{a}\subset R$ be ideals and let $t\in\mathbb{R}_{>0}$.
@@
    If $I\subset J$, $\mathfrak{b}\subset\mathfrak{a}$, and $t\leq s$, then $\overline{I}^{\mathfrak{a}^{t}}\subset\overline{J}^{\mathfrak{b}^{s}}$ and ${}^{\mathfrak{a}^{t}}\overline{I}\subset{}^{\mathfrak{a}^{s}}\overline{J}.$
  4. (d)
    $I^{*\mathfrak{a}^{t}}\subset\overline{I}^{\mathfrak{a}^{t}}$ and ${}^{\mathfrak{a}^{t}}I^{*}\subset{}^{\mathfrak{a}^{t}}\overline{I}.$
@@
    In order to show that $\overline{I}^{\mathfrak{a}^{t}}$ is an ideal, it is enough to check that if $x,y\in\overline{I}^{\mathfrak{a}^{t}},$ then $x+y\in\overline{I}^{\mathfrak{a}^{t}}.$ As $x,y\in\overline{I}^{\mathfrak{a}^{t}},$ there exist $c,d\in R^{\circ}$ such that $cx^{n}\mathfrak{a}^{\lceil tn\rceil}\subset I^{n}$ and $dy^{n}\mathfrak{a}^{\lceil tn\rceil}\subset I^{n}$ for all $n\geq 0.$ Choose $c^{\prime}\in\mathfrak{a}^{2}\cap R^{\circ}.$ Then
    $c^{\prime}cd(x+y)^{n}\mathfrak{a}^{\lceil tn\rceil}=\sum_{j=0}^{n}\binom{n}{j}(cx^{j})(dy^{n-j})c^{\prime}\mathfrak{a}^{\lceil tn\rceil}$ $\subset\sum_{j=0}^{n}\binom{n}{j}(cx^{j})(dy^{n-j})\mathfrak{a}^{\lceil tn\rceil+2}$
    $\subset\sum_{j=0}^{n}\binom{n}{j}(cx^{j}\mathfrak{a}^{\lceil tj\rceil})(dy^{n-j}\mathfrak{a}^{\lceil t(n-j)\rceil})\subset I^{n}.$
    Thus, $\overline{I}^{\mathfrak{a}^{t}}$ is an ideal. Now let $x,y\in{}^{\mathfrak{a}^{t}}\overline{I}.$ Then there exist $c,d\in R^{\circ}$ such that $cx^{n}\mathfrak{a}^{\lceil tn\rceil}\subset\mathfrak{a}^{\lceil tn\rceil}I^{n},$ and $dy^{n}\mathfrak{a}^{\lceil tn\rceil}\subset\mathfrak{a}^{\lceil tn\rceil}I^{n}$ for all $n\geq 0.$ Choose $c^{\prime}\in\mathfrak{a}^{2}\cap R^{\circ}.$ Then
@@
    Since $I\subset J$, $I^{n}\subset J^{n}$ for all $n$, so we have $I^{n}:\mathfrak{a}^{\lceil tn\rceil}\subset J^{n}:\mathfrak{b}^{\lceil sn\rceil}$ for all $n$. Thus, $\overline{I}^{\mathfrak{a}^{t}}\subset\overline{J}^{\mathfrak{b}^{s}}$. Similarly, $I^{n}\mathfrak{a}^{\lceil tn\rceil}:\mathfrak{a}^{\lceil tn\rceil}\subset J^{n}\mathfrak{a}^{\lceil tn\rceil}:\mathfrak{a}^{\lceil tn\rceil}\subset J^{n}\mathfrak{a}^{\lceil sn\rceil}:\mathfrak{a}^{\lceil sn\rceil}$ for all $n$, so ${}^{\mathfrak{a}^{t}}\overline{I}\subset{}^{\mathfrak{a}^{s}}\overline{J}$.
  4. (d)
    Let $x\in I^{*\mathfrak{a}^{t}}.$ Then there exists $c\in R^{\circ}$ such that for all $e\gg 0,$ $cx^{p^{e}}\mathfrak{a}^{\lceil tp^{e}\rceil}\subset I^{[p^{e}]}\subset I^{p^{e}}$ implying that $x\in\overline{I}^{\mathfrak{a}^{t}}.$ One can show the other containment similarly. ∎
@@
Let $I,\mathfrak{a}\subset R$ be ideals and let $t\in\mathbb{R}_{>0}$.
@@
    If $t\geq 1$ then $\overline{I}^{\mathfrak{a}^{t}}\subset\left(\overline{I}^{\mathfrak{a}^{t-1}}:\mathfrak{a}\right)$, and we have equality if $\mathfrak{a}$ is a principal ideal or $t\in\mathbb{Z}$.
@@
  3. (c)
    If $t\in\mathbb{Q}_{>0}$, then $\overline{I}^{\mathfrak{a}^{t}}\subset(\overline{I}:\mathfrak{a}_{t})\subset\overline{I}^{\mathfrak{a}^{\lceil t\rceil}}.$
@@
    Therefore $xy\in\overline{I}^{\mathfrak{a}^{t}}$. Hence $x\in\left(\overline{I}^{\mathfrak{a}^{t-1}}:\mathfrak{a}\right)$.
@@
    Finally, suppose $t\in\mathbb{Z}$ and $x\in\left(\overline{I}^{\mathfrak{a}^{t-1}}:\mathfrak{a}\right)$. Let $y_{1},\ldots,y_{r}$ be the generators of $\mathfrak{a}$. For each $i$, $xy_{i}\in\overline{I}^{\mathfrak{a}^{t-1}}$, and so there exists $c_{i}\in R^{\circ}$ such that for all $n\geq 0$, $c_{i}(xy)^{n}\mathfrak{a}^{(t-1)n}\subset I^{n}$. Let $c=\prod_{i}c_{i}\in R^{\circ}$, and fix $n\geq 0$. All generators of $\mathfrak{a}^{n}$ are of the form $\prod_{i}y_{i}^{n_{i}}$ with $\sum_{i}n_{i}=n$. For such a generator, we have that
@@
    Therefore $cx^{n}\mathfrak{a}^{tn}=cx^{n}\mathfrak{a}^{n}\mathfrak{a}^{(t-1)n}\subset I^{n}$, meaning $x\in\overline{I}^{\mathfrak{a}^{t}}$. Hence $\left(\overline{I}^{\mathfrak{a}^{t-1}}:\mathfrak{a}\right)\subset\overline{I}^{\mathfrak{a}^{t}}$, and we have equality.
@@
    Repeatedly apply the previous part $t$ times.
@@
    Suppose that $x\in(\overline{I}:\mathfrak{a}_{t})$. Let $c\in R^{\circ}$ such that for $n\geq 0$, $c(\overline{I})^{n}\subset I^{n}$. Now, for $n\gg 0$,
    $cx^{n}\mathfrak{a}^{\lceil t\rceil n}=c(x\mathfrak{a}^{\lceil t\rceil})^{n}\subset c(x\mathfrak{a}_{\lceil t\rceil})^{n}\subset c(x\mathfrak{a}_{t})^{n}\subset c(\overline{I})^{n}=I^{n}.$
    Therefore $x\in\overline{I}^{\mathfrak{a}^{\lceil t\rceil}}$. ∎
@@
The reverse containment $(\overline{I}:\mathfrak{a}_{t})\subset\overline{I}^{\mathfrak{a}^{t}}$ in 3.5 is not true in general if $t$ is not an integer. For example, let $R=k[x,y]$, $\mathfrak{a}=(x,y)$, $I=(x^{2},y^{2})$, and $21<t<$ 2. Here $\mathfrak{a}_{t}=(x^{2},xy,y^{2})=\overline{I}$, and so $(\overline{I}:\mathfrak{a}_{t})=R$. However, $1\notin\overline{I}^{\mathfrak{a}^{t}}$. If it were, then there would exist a nonzero $c\in R$ such that for all large $n$, $c\mathfrak{a}^{\lceil tn\rceil}\subset I^{n}$. However, for any $n\in\mathbb{N}$, $\mathfrak{a}^{\lceil tn\rceil}$ is generated by elements of degree $\lceil tn\rceil$, and so $c\mathfrak{a}^{\lceil tn\rceil}$ contains elements of degree $\deg(c)+\lceil tn\rceil$. Since $I$ is generated by elements of degree 2, $I^{n}$ is generated by homogeneous elements of degree $2n$. Thus, we would have that $2n\leq\deg(c)+\lceil tn\rceil$ for all $n$, meaning that $\deg(c)\geq\lceil(2-t)n\rceil$ for all $n$, a contradiction. Therefore $(\overline{I}:\mathfrak{a}_{t})\not\subset\overline{I}^{\mathfrak{a}^{t}}$.
We may use 3.5 to observe that the Hara-Yoshida integral closure is not an idempotent operation on ideals in general.
@@
Let $R$ be a ring, $I\subset R$ an ideal and $f\in R$ a nonzerodivisor. By 3.5, $\overline{I}^{(f)}=(\overline{I}:f)$ and
@@
We conclude this subsection by demonstrating that computation of the $\mathfrak{a}^{t}$-integral closure may be reduced to computation in domains via modding out by minimal primes, unlike Frobenius closure (see 2.3).
@@
Let $\mathfrak{a},I$ be ideals in a reduced ring $R$ and let $t\in\mathbb{R}_{>0}$. An element $x\in\overline{I}^{\mathfrak{a}^{t}}$ if and only if the image of $x$ in $R/\mathfrak{p}$ is in $\overline{((I+\mathfrak{p})/\mathfrak{p})}^{\mathfrak{b}^{t}}$ for all minimal primes $\mathfrak{p}$ of $R.$ Here, $\mathfrak{b}$ denotes the image of $\mathfrak{a}$ in $R/\mathfrak{p}.$
@@
Conversely, assume that for each minimal prime $\mathfrak{p}$ of $R,$ image of $x$ in $R/\mathfrak{p}$ is in $\overline{((I+\mathfrak{p})/\mathfrak{p})}^{\mathfrak{b}^{t}}.$ Let $c_{\mathfrak{p}}\in R\setminus\mathfrak{p}$ such that $c_{\mathfrak{p}}x^{n}\mathfrak{a}^{\lceil tn\rceil}\subset c_{\mathfrak{p}}x^{n}\mathfrak{a}^{\lceil tn\rceil}+\mathfrak{p}\subset I^{n}+\mathfrak{p}$ for all $n.$ By prime avoidance, there exists a non-zero element $d_{\mathfrak{p}}\in R$ which is not in $\mathfrak{p}$ but is in every other minimal prime of $R.$ Then $d_{\mathfrak{p}}c_{\mathfrak{p}}x^{n}\mathfrak{a}^{\lceil tn\rceil}\subset I^{n}+\sqrt{0},$ for all $n.$ Set $c^{\prime}=\sum_{\mathfrak{p}\in\operatorname{MinR}}d_{\mathfrak{p}}c_{\mathfrak{p}}.$ Then $c^{\prime}\in R^{\circ}$ and for all $n$, we have

### v3

@@ 3. The closure operations
In this section we define new operations on ideals whose origin is a formal adaptation of the tight closure adjustments of Hara-Yoshida [HY03] and Vraciu [Vra08] (see 3.11) to integral and Frobenius closures. We want to remind the reader of Remark 2.2, since in this section we use the word closure to refer to both ideal closure operations and pre-closures.
@@ 3.1. The $\mathfrak{a}^{t}$-integral closures
We begin with the adjustments to the ordinary integral closure of an ideal $I$. In this subsection, we consider rings $R$ of arbitrary characteristic, and the ideals $I,\mathfrak{a}\subset R$ we consider will typically be assumed to be of positive height, so that they intersect $R^{\circ}$.
@@
Let $I,\mathfrak{a}\subset R$ be ideals and let $t\in\mathbb{R}_{>0}$. Then, we define the Hara-Yoshida $\mathfrak{a}^{t}$-Integral closure of $I$, denoted $\overline{I}^{\mathfrak{a}^{t}}$, by
[…]
Similarly, we define the Vraciu $\mathfrak{a}^{t}$-Integral closure of $I$, denoted $^{\mathfrak{a}^{t}}\overline{I}$, by
[…]
We will later show that the Hara-Yoshida $\mathfrak{a}^{t}$-integral closure is a pre-closure (3.3) and not an ideal closure operation (3.7). Further, in virtue of being equal to the classical integral closure (3.3), the Vraciu $\mathfrak{a}^{t}$-integral closure is an ideal closure operation.
We first note that the “for all $n\gg 0$” in the definitions above may be strengthened to “for all nn”, so long as $I$ is of positive height.
@@
Let $R$ be a ring and $I,\mathfrak{a}\subset R$ be ideals where $I$ has positive height, and let $t\in\mathbb{R}_{>0}$. Then $x\in\overline{I}^{\mathfrak{a}^{t}}$ (respectively $x\in{}^{\mathfrak{a}^{t}}\overline{I}$) if and only if there exists $c\in R^{\circ}$ such that, for all $n\geq 0$, $cx^{n}\mathfrak{a}^{\lceil tn\rceil}\subset I^{n}$ (respectively $cx^{n}\mathfrak{a}^{\lceil tn\rceil}\subset I^{n}\mathfrak{a}^{\lceil tn\rceil}$).
@@
The proofs of both parts are similar, so we provide the proof for the Hara-Yoshida version only. Let $x\in\overline{I}^{\mathfrak{a}^{t}}.$ Then there exist $n_{0}\in\mathbb{N}$ and $c\in R^{\circ}$ such that $cx^{n}\mathfrak{a}^{\lceil tn\rceil}\subset I^{n}$ for all $n\geq n_{0}.$ Choose $b\in R^{\circ}\cap I^{n_{0}}$. Then $bc\in R^{\circ}$ and $bcx^{n}\mathfrak{a}^{\lceil tn\rceil}\subset I^{n}$ for all $n\geq 0$. ∎
We now record several basic properties of these closure operations.
@@
Let $R$ be a ring, $I,\mathfrak{a}\subset R$ be ideals of positive height, and $t\in\mathbb{R}_{>0}$.
@@
    If $I\subset J$, $\mathfrak{b}\subset\mathfrak{a}$, and $t\leq s$, then $\overline{I}^{\mathfrak{a}^{t}}\subset\overline{J}^{\mathfrak{b}^{s}}$.
@@
    In order to show that $\overline{I}^{\mathfrak{a}^{t}}$ is an ideal, it is enough to check that if $x,y\in\overline{I}^{\mathfrak{a}^{t}},$ then $x+y\in\overline{I}^{\mathfrak{a}^{t}}.$ As $x,y\in\overline{I}^{\mathfrak{a}^{t}},$ there exist $c,d\in R^{\circ}$ such that $cx^{n}\mathfrak{a}^{\lceil tn\rceil}\subset I^{n}$ and $dy^{n}\mathfrak{a}^{\lceil tn\rceil}\subset I^{n}$ for all $n\geq 0.$ Choose $c^{\prime}\in\mathfrak{a}\cap R^{\circ}.$ Then
    $c^{\prime}cd(x+y)^{n}\mathfrak{a}^{\lceil tn\rceil}=\sum_{j=0}^{n}\binom{n}{j}(cx^{j})(dy^{n-j})c^{\prime}\mathfrak{a}^{\lceil tn\rceil}$ $\subset\sum_{j=0}^{n}\binom{n}{j}(cx^{j})(dy^{n-j})\mathfrak{a}^{\lceil tn\rceil+1}$
    $\subset\sum_{j=0}^{n}\binom{n}{j}(cx^{j}\mathfrak{a}^{\lceil tj\rceil})(dy^{n-j}\mathfrak{a}^{\lceil t(n-j)\rceil})\subset I^{n}$
    since for any $j$, $\lceil tn\rceil+1\geq\lceil tj\rceil+\lceil t(n-j)\rceil$ Thus, $\overline{I}^{\mathfrak{a}^{t}}$ is an ideal. Similarly, let $x,y\in{}^{\mathfrak{a}^{t}}\overline{I}.$ Then there exist $c,d\in R^{\circ}$ such that $cx^{n}\mathfrak{a}^{\lceil tn\rceil}\subset\mathfrak{a}^{\lceil tn\rceil}I^{n},$ and $dy^{n}\mathfrak{a}^{\lceil tn\rceil}\subset\mathfrak{a}^{\lceil tn\rceil}I^{n}$ for all $n\geq 0.$ Choose $c^{\prime}\in\mathfrak{a}\cap R^{\circ}.$ Then
@@
    Since $I\subset J$, $I^{n}\subset J^{n}$ for all $n$, so we have $I^{n}:\mathfrak{a}^{\lceil tn\rceil}\subset J^{n}:\mathfrak{b}^{\lceil sn\rceil}$ for all $n$. Thus, $\overline{I}^{\mathfrak{a}^{t}}\subset\overline{J}^{\mathfrak{b}^{s}}$. ∎
@@
Let $R$ be a ring, $I,\mathfrak{a}\subset R$ be ideals of positive height and $t\in\mathbb{R}_{>0}$.
@@
    If $t>1$ then $\overline{I}^{\mathfrak{a}^{t}}\subset\left(\overline{I}^{\mathfrak{a}^{t-1}}:\mathfrak{a}\right)$, and we have equality if $\mathfrak{a}$ is a principal ideal or $t\in\mathbb{Z}$.
@@
    If $t\in\mathbb{Q}_{>0}$, then $\overline{I}^{\mathfrak{a}^{t}}\subset(\overline{I}:\mathfrak{a}_{t})\subset\overline{I}^{\mathfrak{a}^{\lceil t\rceil}},$ where $\mathfrak{a}_{t}$ is the $t$-th rational power of $\mathfrak{a}$.
  3. (c)
@@
    Therefore $xy\in\overline{I}^{\mathfrak{a}^{t-1}}$. Hence $x\in\left(\overline{I}^{\mathfrak{a}^{t-1}}:\mathfrak{a}\right)$.
@@
    Finally, suppose $t\in\mathbb{Z}$ and $x\in\left(\overline{I}^{\mathfrak{a}^{t-1}}:\mathfrak{a}\right)$. Let $y_{1},\ldots,y_{r}$ be the generators of $\mathfrak{a}$. For each $i$, $xy_{i}\in\overline{I}^{\mathfrak{a}^{t-1}}$, and so there exists $c_{i}\in R^{\circ}$ such that for all $n\geq 0$, $c_{i}(xy_{i})^{n}\mathfrak{a}^{(t-1)n}\subset I^{n}$. Let $c=\prod_{i}c_{i}\in R^{\circ}$, and fix $n\geq 0$. All generators of $\mathfrak{a}^{n}$ are of the form $\prod_{i}y_{i}^{n_{i}}$ with $\sum_{i}n_{i}=n$. For such a generator, we have that
@@
    Therefore $cx^{n}\mathfrak{a}^{tn}=cx^{n}\mathfrak{a}^{n}\mathfrak{a}^{(t-1)n}\subset I^{n}$, meaning $x\in\overline{I}^{\mathfrak{a}^{t}}$. Hence $\left(\overline{I}^{\mathfrak{a}^{t-1}}:\mathfrak{a}\right)\subset\overline{I}^{\mathfrak{a}^{t}}$.
@@
    Suppose that $x\in(\overline{I}:\mathfrak{a}_{t})$. Let $c\in R^{\circ}$ such that for $n\gg 0$, $c(\overline{I})^{n}\subset I^{n}$ ([SH06, Corollary 6.8.12]). Now, for $n\gg 0$,
    $cx^{n}\mathfrak{a}^{\lceil t\rceil n}=c(x\mathfrak{a}^{\lceil t\rceil})^{n}\subset c(x\mathfrak{a}_{\lceil t\rceil})^{n}\subset c(x\mathfrak{a}_{t})^{n}\subset c(\overline{I})^{n}\subset I^{n}.$
    Therefore $x\in\overline{I}^{\mathfrak{a}^{\lceil t\rceil}}$.
@@
    Applying part (b) with $t\in\mathbb{Z}_{\geq 1}$, we have that $\overline{I}^{\mathfrak{a}^{t}}=(\overline{I}:\mathfrak{a}_{t})=(\overline{I}:\overline{\mathfrak{a}^{t}})=(\overline{I}:\mathfrak{a}^{t}).$
@@
The reverse containment $(\overline{I}:\mathfrak{a}_{t})\subset\overline{I}^{\mathfrak{a}^{t}}$ in 3.5 is not true in general if $t$ is not an integer. For example, let $R=k[x,y]$, $\mathfrak{a}=(x,y)$, $I=(x^{2},y^{2})$, and $1<t<2$. Since $\mathfrak{a}$ is generated by degree 1 elements, $\mathfrak{a}_{t}$ is generated by elements of degree at least $t$, and so $\mathfrak{a}_{t}\subset(x^{2},xy,y^{2})$. On the other hand, $\mathfrak{a}_{t}\supseteq\mathfrak{a}_{2}=\overline{\mathfrak{a}^{2}}=(x^{2},xy,y^{2})$, and so we have equality. Therefore $\mathfrak{a}_{t}=(x^{2},xy,y^{2})=\overline{I}$, and so $(\overline{I}:\mathfrak{a}_{t})=R$. However, $1\notin\overline{I}^{\mathfrak{a}^{t}}$. If it were, then there would exist a nonzero $c\in R$ such that for all large $n$, $c\mathfrak{a}^{\lceil tn\rceil}\subset I^{n}$. However, for any $n\in\mathbb{N}$, $\mathfrak{a}^{\lceil tn\rceil}$ is generated by elements of degree $\lceil tn\rceil$, and so $c\mathfrak{a}^{\lceil tn\rceil}$ contains elements of degree $\deg(c)+\lceil tn\rceil$. Since $I$ is generated by elements of degree 2, $I^{n}$ is generated by homogeneous elements of degree $2n$. Thus, we would have that $2n\leq\deg(c)+\lceil tn\rceil$ for all large $n$, meaning that $\deg(c)\geq\lceil(2-t)n\rceil$ for all large $n$, a contradiction. Therefore $(\overline{I}:\mathfrak{a}_{t})\not\subset\overline{I}^{\mathfrak{a}^{t}}$.
We may use 3.5 to observe that the $\mathfrak{a}^{t}$-integral closure is not an idempotent operation on ideals in general.
@@
Let $R$ be a ring, $I\subset R$ an ideal of positive height and $f\in R$ a nonzerodivisor. By 3.5, $\overline{I}^{(f)}=(\overline{I}:f)$ and
@@
We now demonstrate that the computation of the $\mathfrak{a}^{t}$-integral closure may be reduced to a computation in domains via going modulo the minimal primes, like the usual tight and integral closures.
@@
Let $\mathfrak{a},I$ be ideals of positive height in a reduced ring $R$, $t\in\mathbb{R}_{>0}$, and $x\in R$. Then $x\in\overline{I}^{\mathfrak{a}^{t}}$ if and only if the image of $x$ in $R/\mathfrak{p}$ is in $\overline{((I+\mathfrak{p})/\mathfrak{p})}^{\mathfrak{b}^{t}}$ for all minimal primes $\mathfrak{p}$ of $R$, where $\mathfrak{b}$ denotes the image of $\mathfrak{a}$ in $R/\mathfrak{p}.$
@@
Conversely, assume that for each minimal prime $\mathfrak{p}$ of $R,$ the image of $x$ in $R/\mathfrak{p}$ is in $\overline{((I+\mathfrak{p})/\mathfrak{p})}^{\mathfrak{b}^{t}}.$ Let $c_{\mathfrak{p}}\in R\setminus\mathfrak{p}$ such that $c_{\mathfrak{p}}x^{n}\mathfrak{a}^{\lceil tn\rceil}\subset c_{\mathfrak{p}}x^{n}\mathfrak{a}^{\lceil tn\rceil}+\mathfrak{p}\subset I^{n}+\mathfrak{p}$ for all $n.$ By prime avoidance, there exists a non-zero element $d_{\mathfrak{p}}\in R$ which is not in $\mathfrak{p}$ but is in every other minimal prime of $R.$ Then $d_{\mathfrak{p}}c_{\mathfrak{p}}x^{n}\mathfrak{a}^{\lceil tn\rceil}\subset I^{n}+\sqrt{0},$ for all $n.$ Set $c^{\prime}=\sum_{\mathfrak{p}\in\operatorname{MinR}}d_{\mathfrak{p}}c_{\mathfrak{p}}.$ Then $c^{\prime}\in R^{\circ}$ and for all $n$, we have